OpenStudy (calculusxy):

MEDAL!!! Find the correct solution that best describes the inequality. -3 < \(\large \frac{5x+9}{7}\) ≤ 2 a) (-∞, 1) ∩ (-6,∞) b) (-∞, 1] ∩ (-6, ∞) c) (-∞, 1) ∩ [-6, ∞) d) (-∞, 1] ∩ [-6, ∞)

OpenStudy (calculusxy):

@Nnesha @Hero @SolomonZelman

OpenStudy (calculusxy):

@jim_thompson5910

OpenStudy (calculusxy):

I tried to solve for this and got that \(\large -6 < x \le 1\)

jimthompson5910 (jim_thompson5910):

hint: break up the inequality to get -3 < (5x+9)/7 and (5x+9)/7 <= 2

OpenStudy (calculusxy):

Would \(\color{red}{\large 6 < x \le -1}\) be correct?

jimthompson5910 (jim_thompson5910):

yes but you need to find an equivalent form

OpenStudy (calculusxy):

I don't understand

jimthompson5910 (jim_thompson5910):

actually no, sorry, the answer is -6 < x <= 1 you need to find an equivalent form to that

jimthompson5910 (jim_thompson5910):

solving `-3 < (5x+9)/7` leads to what?

OpenStudy (calculusxy):

-6 < x

jimthompson5910 (jim_thompson5910):

write that in interval notation

OpenStudy (calculusxy):

(-6, \(\infty\))

jimthompson5910 (jim_thompson5910):

good

OpenStudy (calculusxy):

But I noticed that when I drew the two inequalities together they did overlap.